Please help I am a little hung up on this one.
Find the length of x in a triangle. Express yourself in simplified radical form.
Left side of triangle: 10
Rigth side of triangle: x
Bottom of triangle: 8

Answer by stanbon(75887)

Find the length of x in a triangle. Express yourself in simplified radical form.
Left side of triangle: 10
Rigth side of triangle: x
Bottom of triangle: 8
-------------------------
Since the left side is the hypotenuse:
10^2 = x^2 + 8^2
x^2 = 100-64
x^2 = 36
x = sqrt(36)
x = 6
===========
Cheers,
Stan H.